6-11b(less than or equal to)-3(negative 3)
Answer by checkley77(12844)

6-11b<=-3
-11b<=-3-6
-11b<=-9
b>=-9/-11 (note: the < sign changes to a > sign when dividing by a negative value).
b>=9/11 answer.
Proof:
Let b=10/11
6-11*10/11<=-3
6-10<=-3
-4<=-3